OpenStudy (anonymous):

which is a simplified form of the expression 2c+2+6c-4

OpenStudy (anonymous):

Simplifying 2c + 2 + 6c + -4 = 0 Reorder the terms: 2 + -4 + 2c + 6c = 0 Combine like terms: 2 + -4 = -2 -2 + 2c + 6c = 0 Combine like terms: 2c + 6c = 8c -2 + 8c = 0 Solving -2 + 8c = 0 Solving for variable 'c'. Move all terms containing c to the left, all other terms to the right. Add '2' to each side of the equation. -2 + 2 + 8c = 0 + 2 Combine like terms: -2 + 2 = 0 0 + 8c = 0 + 2 8c = 0 + 2 Combine like terms: 0 + 2 = 2 8c = 2 Divide each side by '8'. c = 0.25 Simplifying c = 0.25